's age is 7/10 onicolef joseph's age. In 4 years, nicole's age will be 8/11 of Joseph's age. How old is nicole now?
Answer by Cromlix(4381)

nicole's age is 7/10 of joseph's age.
N = 7/10J
In 4 years, nicole's age will be 8/11 of Joseph's age.
N + 4 = 8/11(J + 4)
Replace N with 7/10J
7/10J + 4 = 8/11J + 32/11
7/10J - 8/11J = 32/11 - 44/11 (4 = 44/11)
-3/110J = -12/11
J = (-12/11)/(-3/110)
J = -12 * 110/11 * -3
J = 40
Nicole's current age is 32 years old.
Hope this helps.
:-)